# Effective and Efficient Conversation Retrieval for Dialogue State Tracking with Implicit Text Summaries

### v2

@@ Abstract
Few-shot dialogue state tracking (DST) with Large Language Models (LLM) relies on an effective and efficient conversation retriever to find similar in-context examples for prompt learning. Previous works use raw dialogue context as search keys and queries, and a retriever is fine-tuned with annotated dialogues to achieve superior performance. However, the approach is less suited for scaling to new domains or new annotation languages, where fine-tuning data is unavailable. To address this problem, we handle the task of conversation retrieval based on text summaries of the conversations. A LLM-based conversation summarizer is adopted for query and key generation, which enables effective maximum inner product search. To avoid the extra inference cost brought by LLM-based conversation summarization, we further distill a light-weight conversation encoder which produces query embeddings without decoding summaries for test conversations. We validate our retrieval approach on MultiWOZ datasets with GPT-Neo-2.7B and LLaMA-7B/30B. The experimental results show a significant improvement over relevant baselines in real few-shot DST settings.
@@ 2 Conversation Retrieval with Summaries for LLM-based DST
There are two major challenges of conversation retrieval for LLM-based DST described above. First, a good representation of search keys and queries need to be found. As we analyze in Section 1, the similarity of two dialogues is not directly quantifiable by semantic distance, but rather requires more sophisticated structural matching mechanism or a higher-order similarity function. This requirement leads to the second challenge as to how to train an effective conversation retriever that can scale across domains. Previous works (Hu et al., 2022; Chen et al., 2023) mainly fine-tune pre-trained dense retrievers with annotated dialogues obtained from the support set. However, fine-tuning is not realistic in a real few-shot setting and for every domain.
@@
Figure 1: Comparison btw. (a) off-the-shelf retriever with query generation and (b) CONVERSE w/o query generation.
@@ 3.1 Model
In our problem setup, we are given a set of unlabeled conversations between a user and system, denoted as $\mathcal{D}^{u}=\{{\mathbf{x}}_{i}\}_{i=1}^{n}$ where each conversation ${\mathbf{x}}_{i}$ consists of $l_{i}$ utterances $({\mathbf{u}}_{i,j},\ldots,{\mathbf{u}}_{i,l_{i}})$ and each utterance ${\mathbf{u}}_{i,j}$ is a sequence of $T_{i,j}$ tokens $(x_{i,j,1},\ldots,x_{i,j,T_{i,j}})$. As shown in Figure 2-(b), the training data of CONVERSE is prepared by invoking the conversation summarizer to generate a summary for each conversation ${\mathbf{x}}_{i}$, denoted as ${\mathbf{z}}_{i}$, which consists of $T^{\prime}_{i}$ tokens with $T^{\prime}_{i}\ll\sum_{j=1}^{l_{i}}T_{i,j}$. We denote the dataset augmented with summaries as $\mathcal{D}^{a}=\{({\mathbf{x}}_{i},{\mathbf{z}}_{i})\}_{i=1}^{n}$. For brevity, we omit the first subscript $i$ if there is no ambiguity. Given the set of conversation-summary pairs, the goal is to train an encoder $f_{\theta}:\mathbb{N}^{T}\to\mathbb{R}^{T\times d}$ such that the similarity between a conversation and its summary is maximized.
@@
$\hat{f}_{\theta,\phi}({\mathbf{x}})=\begin{bmatrix}w_{1,1}f_{\theta}({\mathbf{x}})^{\top}_{1,1}\\ \vdots\\ w_{l-1,T_{l-1}}f_{\theta}({\mathbf{x}})^{\top}_{l-1,T_{l-1}}\\ f_{\theta}({\mathbf{x}})^{\top}_{l,1}\\ \vdots\\ f_{\theta}({\mathbf{x}})^{\top}_{l,T_{l}}\end{bmatrix}\in\mathbb{R}^{T\times d}$ (7)
@@ 4.1 Experimental Setup
    IC-DST (Hu et al., 2022): It utilizes dialogue labels to construct positive and negative pairs for fine-tuning a pretrained SBERT (Reimers and Gurevych, 2019). The retrieval key is a dialogue context, and the best dialogue context is reported to be previous dialogue state + current user input (which is better than a full dialogue).
@@
Second, we evaluate the distilled conversation encoder (CONVERSE). To train CONVERSE, we use the same conversation summarizer to generate a summary for every turn of every conversation from the full MultiWOZ training set, resulting in a total of 56,776 conversation-summary pairs. The parameters $\theta$ of the dual encoder $f_{\theta}$ and $\hat{f}_{\theta,\phi}$ are shared and initialized with LinkBERT (Yasunaga et al., 2022), and trained on the conversation-summary pairs for 20 epochs with the objective in equation 9. LinkBERT (Yasunaga et al., 2022) is chosen since we empirically find that it offers the best general-purpose weight of initialization. We use the AdamW optimizer (Loshchilov and Hutter, 2018) with learning rate $5\cdot 10^{-5}$ and batch size 200. We use 8 A100 GPUs for training the model.
@@ 4.2 Quantitative Results
Table 2: Performance of LLaMA-30B w/ 100 labeled conv.
The DST results are shown in Table 1 and Table 2. The first set of comparisons is between conversation retrieval with and without explicit query/key generation. We observe that using the summary as search keys/queries significantly improves the E2E results, when evaluated with the same off-the-shelf retriever (GTR-T5 or Jina). The result is slightly behind IC-DST which fine-tunes the retriever with dialogue state information in the key. However, after introducing the distilled CONVERSE model, we achieve much better E2E results than all baselines. Although our motivation of distilling a conversation encoder is to reduce the inference cost, it turns out that the light-weight model is also helpful in E2E performance. We hypothesize that the improved performance brought by CONVERSE is attributed to two factors. The first and foremost is that we leverage a dual encoder architecture to optimize the matching between conversation-summary pairs. This suggests that the retrieval component is optimized for the task-specific keys and values. A secondary explanation of the performance gain is that the conversation encoder avoids error propagation in explicit summary decoding and re-embedding. It should be noted that the above findings are consistent across different datasets (MultiWOZ 2.1/2.4) and language models (GPT-Neo and LLaMA-7B/30B).
@@ 4.4 Qualitative Results
Examples of the generated summaries are shown in Table 4 and  5. For the first example, the model generates the summary about booking a taxi. It is noteworthy that the model focuses on the latest user utterance while disregarding previous user requests for hotel and restaurant reservation. For the second example, the model misses out on the arrival time for generating the summary. Identification and correction of such errors are topics we will explore in future work. We include more examples in Appendix 12.
@@ Acknowledgement
We would like to express our sincere gratitude to Dominik Wagner for the numerous helpful pieces of feedbacks.
@@ Appendix C Example of Summaries
In Tables 8 to 12, we provide more summaries generated by the LLM, gpt-3.5-turbo.

### v3

@@ Abstract
Few-shot dialogue state tracking (DST) with Large Language Models (LLM) relies on an effective and efficient conversation retriever to find similar in-context examples for prompt learning. Previous works use raw dialogue context as search keys and queries, and a retriever is fine-tuned with annotated dialogues to achieve superior performance. However, the approach is less suited for scaling to new domains or new annotation languages, where fine-tuning data is unavailable. To address this problem, we handle the task of conversation retrieval based on text summaries of the conversations. A LLM-based conversation summarizer is adopted for query and key generation, which enables effective maximum inner product search. To avoid the extra inference cost brought by LLM-based conversation summarization, we further distill a light-weight conversation encoder which produces query embeddings without decoding summaries for test conversations. We validate our retrieval approach on MultiWOZ datasets with GPT-Neo-2.7B and LLaMA-7B/30B. The experimental results show a significant improvement over relevant baselines in few-shot DST settings.
@@ 2 Conversation Retrieval with Summaries for LLM-based DST
There are two major challenges of conversation retrieval for LLM-based DST described above. First, a good representation of search keys and queries need to be found. As we analyze in Section 1, the similarity of two dialogues is not directly quantifiable by semantic distance, but rather requires more sophisticated structural matching mechanism or a higher-order similarity function. This requirement leads to the second challenge as to how to train an effective conversation retriever that can scale across domains. Previous works (Hu et al., 2022; Chen et al., 2023) mainly fine-tune pre-trained dense retrievers with annotated dialogues obtained from the support set. However, fine-tuning is not realistic in a few-shot setting and for every domain.
@@
Figure 1: Comparison between (a) off-the-shelf retriever with query generation and (b) CONVERSE w/o query generation.
@@ 3.1 Model
In our problem setup, we are given a set of unlabeled conversations between a user and system, denoted as $\mathcal{D}^{u}=\{{\mathbf{x}}_{i}\}_{i=1}^{n}$ where each conversation ${\mathbf{x}}_{i}$ consists of $l_{i}$ utterances $({\mathbf{u}}_{i,j},\ldots,{\mathbf{u}}_{i,l_{i}})$ and each utterance ${\mathbf{u}}_{i,j}$ is a sequence of $T_{i,j}$ tokens $(x_{i,j,1},\ldots,x_{i,j,T_{i,j}})$. As shown in Figure 2-(b), the training data of CONVERSE is prepared by invoking the conversation summarizer to generate a summary for each conversation ${\mathbf{x}}_{i}$, denoted as ${\mathbf{z}}_{i}$, which consists of $T^{\prime}_{i}$ tokens with $T^{\prime}_{i}\ll\sum_{j=1}^{l_{i}}T_{i,j}$. We denote the dataset augmented with summaries as $\mathcal{D}^{a}=\{({\mathbf{x}}_{i},{\mathbf{z}}_{i})\}_{i=1}^{n}$. For brevity, we omit the first subscript $i$ if there is no ambiguity. Given the set of conversation-summary pairs, the goal is to train an encoder $f_{\theta}:\mathcal{V}^{T}\to\mathbb{R}^{T\times d}$ such that the similarity between a conversation and its summary is maximized, where $\mathcal{V}$ denotes a set of predefined tokens.
@@
$\hat{f}_{\theta,\phi}({\mathbf{x}})=\begin{bmatrix}w_{1,1}f_{\theta}({\mathbf{x}})^{\top}_{1,1}\\ \vdots\\ w_{l-1,T_{l-1}}f_{\theta}({\mathbf{x}})^{\top}_{l-1,T_{l-1}}\\ f_{\theta}({\mathbf{x}})^{\top}_{l,1}\\ \vdots\\ f_{\theta}({\mathbf{x}})^{\top}_{l,T_{l}}\end{bmatrix}\in\mathbb{R}^{T\times d}$
@@ 4.1 Experimental Setup
    IC-DST (Hu et al., 2022): It utilizes dialogue labels to construct positive and negative pairs for fine-tuning a pretrained SBERT (Reimers and Gurevych, 2019) or LinkBERT (Yasunaga et al., 2022) as a retriever. The retrieval key is a dialogue context, and the best dialogue context is reported to be previous dialogue state + current user input (which is better than a full dialogue).
@@
Second, we evaluate the distilled conversation encoder (CONVERSE). To train CONVERSE, we use the same conversation summarizer to generate a summary for every turn of every conversation from the full MultiWOZ training set, resulting in a total of 56,776 conversation-summary pairs. The parameters $\theta$ of the dual encoder $f_{\theta}$ and $\hat{f}_{\theta,\phi}$ are shared and initialized with LinkBERT (Yasunaga et al., 2022), and trained on the conversation-summary pairs for 20 epochs with the objective in equation 2. LinkBERT (Yasunaga et al., 2022) is chosen since we empirically find that it offers the best general-purpose weight of initialization. We use the AdamW optimizer (Loshchilov and Hutter, 2018) with learning rate $5\cdot 10^{-5}$ and batch size 200. We use eight A100 GPUs for training the model.
@@ 4.2 Quantitative Results
Table 2: JGA and F1 of LLaMA-30B with 100 labeled conversations.
Main Results
The DST results are shown in Table 1 and Table 2. The first set of comparisons is between conversation retrieval with and without explicit query/key generation. We observe that using the summary as search keys/queries significantly improves the end-to-end (E2E) results, when evaluated with the same off-the-shelf retriever (GTR-T5 or Jina). The result is slightly behind IC-DST which fine-tunes the retriever with dialogue state information in the key. However, after introducing the distilled CONVERSE model, we achieve much better E2E results than all baselines. Although our motivation of distilling a conversation encoder is to reduce the inference cost, it turns out that the light-weight model is also helpful in E2E performance. We hypothesize that the improved performance brought by CONVERSE is attributed to two factors. The first and foremost is that we leverage a dual encoder architecture to optimize the matching between conversation-summary pairs. This suggests that the retrieval component is optimized for the task-specific keys and values. A secondary explanation of the performance gain is that the conversation encoder avoids error propagation in explicit summary decoding and re-embedding. It should be noted that the above findings are consistent across different datasets (MultiWOZ 2.1/2.4) and language models (GPT-Neo and LLaMA-7B/30B).
JGA
Model MWZ-2.1 MWZ-2.4
DS2 + BART-Large $7.60$±$\pm 2.17$ $5.86$±$\pm 4.52$
DS2 + T5-Large $17.71$±$\pm 1.84$ $19.08$±$\pm 1.23$
CONVERSE + LLaMA-7B 19.3319.33{19.33}19.33±$\pm 0.91$ 20.3520.35{20.35}20.35±$\pm 1.03$
CONVERSE + LLaMA-30B 27.35±$\pm 0.77$ 28.23±$\pm 1.58$
Table 3: Comparison against few-shot finetuning methods.
Comparison Against Few-Shot Finetuning
Recently, Mosbach et al. (2023) have shown that few-shot fine-tuning outperforms in-context learning in some settings, which makes people wonder how few-shot fine-tuning behaves with 100 labeled dialogues for dialogue state tracking tasks. To answer this question, we compare our method, CONVERSE, against one of the strongest few-shot fine-tuning methods, DS2 (Shin et al., 2022), using BART (Lewis et al., 2020) and T5 (Raffel et al., 2020) language models. As shown in Table 3, our method, CONVERSE, outperforms the few-shot fine-tuning method, DST. Note that the BART-based DS2 severely overfits to the small labeled dataset and the T5-based model performs worse than the in-context learning method, even though T5 model is pretrained on an additional large-scale labeled dialogue summarization dataset, SAMSum (Gliwa et al., 2019).
MultiWOZ 2.1 MultiWOZ 2.4
Model JGA F1 JGA F1
LLaMA-7B (Touvron et al., 2023)
IC-DST (SBERT) $12.52$±$\pm 0.68$ $62.11$±$\pm 0.38$ $12.43$±$\pm 0.09$ $62.45$±$\pm 0.87$
CONVERSE 14.05±$\pm 0.58$ 63.37±$\pm 1.53$ 14.23±$\pm 0.48$ 64.18±$\pm 1.47$
Table 4: Out-of domain generalization using 100 labeled conversations with LLaMA-7B.
Out-of Domain Generalization
To verify our hypothesis that our unsupervised retriever CONVERSE generalizes better to unseen domain than supervised methods, we hold out the hotel domain from the MultiWOZ dataset and train the retrievers, IC-DST and CONVERSE on the remaining four domains: train, restaurant, taxi, and attraction. Then we evaluate the performance of the few-shot in-context learning with the retrievers on test examples from the unseen domain, hotel. As shown in Table 4, our model CONVERSE outperforms IC-DST by a large margin, which empirically validates that our unsupervised retriever generalizes better to unseen domain than the supervised one.
@@ 4.4 Qualitative Results
Examples of the generated summaries are shown in Table 6 and  7. For the first example, the model generates the summary about booking a taxi. It is noteworthy that the model focuses on the latest user utterance while disregarding previous user requests for hotel and restaurant reservation. For the second example, the model misses out on the arrival time for generating the summary. Identification and correction of such errors are topics we will explore in future work. We include more examples in Appendix 14.
@@ Acknowledgement
We would like to express our sincere gratitude to Dominik Wagner and Minki Kang for the numerous helpful pieces of feedbacks.
@@ Appendix C Example of Summaries
In Tables 10 to 14, we provide more summaries generated by the LLM, gpt-3.5-turbo.